OpenStudy (anonymous):

The sum of two numbers is 44 and the difference is 10. What are the numbers?

OpenStudy (cp9454):

let two numbers be a and b, a+b= 44 a-b= 10, solve the eqn and you will get value of a and b

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

ok since the 2 numbers add to get 44, that is represented by a+b=44 but when subtracted they give 10 so that is represented by a-b=10 we can find a by solving for b in the first equation and then plugging it into the other one. to solve for be we subtract a from both sides and get b=44-a. so if we plug that into the other equations we get a-(44-a)=10 a-44+a=10 2a-44=10 now add 44 to both sides 2a=54 now divide by 2 a=27 now we have found one of the numbers

OpenStudy (anonymous):

To find b, or the other number, it is much easier. 27+b=44 subtract 27 from both sides to find the other one!

OpenStudy (anonymous):

got it!! Thank you :)